In order to solve 7=n+5/3 you have to do this N+5/3=7 and then you make N+5/3*7/1 and you cross multiply to get N+5=21 and then you subtract 5 from both sides and you get N=16.
Found 2 solutions by richard1234, Intellectualkid1337:
Answer by richard1234(7193)

Do you mean

, or

?

Answer by Intellectualkid1337(1)

In order to solve 7=n+5/3 you have to do this N+5/3=7 and then you make N+5/3*7/1 and you cross multiply to get N+5=21 and then you subtract 5 from both sides and you get N=16.